Which label describes the following?
3x + 6 = 10
Expression
Linear equation
Equation which is not linear

Answer by praseenakos@yahoo.com(507)

QUESTION:
Which label describes the following?
3x + 6 = 10
Expression
Linear equation
Equation which is not linear

ANSWER:
EXPLANATION"

A linear equation is an equation involving only the sum of constants or products of constants and the first power of a variable.

Such an equation is equivalent to equating a first-degree polynomial to zero.

These equations are called "linear" because they represent straight lines in Cartesian coordinates.

A common form of a linear equation in two variables is y = mx + b,

(e.g. y = 3x + 5).

Equations involving terms such as x2, y1/3, and xy are "non-linear".

================================================================================
An expression is a combination of numbers, operators, grouping symbols (such as brackets and parentheses) and/or free variables and bound variables arranged in a meaningful way which can be evaluated.

For a given combination of values for the free variables, an expression may be evaluated, although for some combinations of values of the free variables, the expression may be undefined. Thus an expression represents a function whose inputs are the values assigned the free variables and whose output is the resulting value of the expression.
For example, the expression
x / y
evaluated for x = 10, y = 5, outputs the number 2; but is undefined for y = 0
================================================================================
Now let's come back to your Question:


3x + 6 = 10

This is an expression.
This is linear equation.


Hope you understood.

Regards.

praseenakos@yahoo.co.in